1. For what values of h the vectors


⟶ [ 1 ⟶ [ -1 ⟶ [ 5

u = -3 u2 = 9 u3 = -7 <--- u1, u2, and u3 are vectors

-2] -6] h]


are linearly independent?


1
Expert's answer
2022-03-21T02:39:31-0400

"\vec{x_1}=(1,-1,5), \vec{x_2}=(-3,9,-7), \vec{x_3}=(-2,-6,h)\\
k_1\vec{x_1}+k_2\vec{x_2}+k_3\vec{x_3}=\vec{0}\\
k_1=k_2=k_3=0\\
k_1(1,-1,5)+k_2(-3,9,-7)+k_3(-2,-6,h)=\\=(0,0,0)\\
k_1-3k_2-2k_3=0\\
-k_1+9k_2-6k_3=0\\
5k_1-7k_2+hk_3=0\\
\Delta=
\begin{vmatrix}
 1& -3&-2 \\
 -1 & 9&-6\\
5&-7&h
\end{vmatrix}=9h+90-14+\\+90-3h-42=6h+124\neq0\\
h\neq-\frac{124}{6}\\
h\neq-\frac{62}{3}\\"